# Uncertainty Quantification in Retrieval Augmented Question Answering

### v1

@@ 3.2 Answer Uncertainty Estimation
For our QA task, we want to define an estimator $\{x,R\}\mapsto\text{{u}}_{\mathcal{M}}(\{x,R\})$ which quantifies the uncertainty of model $\mathcal{M}$ when generating answer $y$ for question $x$ based on a prompt with passages $R$. We propose estimating $\text{{u}}_{\mathcal{M}}$ directly from the utility scores of individual passages in $R$. The key intuition is that the higher the utility of one (or more) passages, the less uncertain $\mathcal{M}$ will be when generating answer $y$. Specifically, we estimate answer uncertainty $\textbf{u}_{\mathcal{M}}$ by taking the maximum utility score among the passages in $R$:
@@ 4 Experimental Setup
We evaluate our approach to predicting answer uncertainty on short-form question answering tasks. Specifically, we present experiments on the following six datasets: Natural Questions Kwiatkowski et al. (2019), TriviaQA Joshi et al. (2017), WebQuestions Berant et al. (2013), SQuAD (Rajpurkar et al., 2016), and PopQA Mallen et al. (2023). We also evaluate on RefuNQ Liu et al. (2024a), a dataset with unanswerable questions about non-existing entities. In Appendix B.1, we describe each dataset, provide example questions, and make available details about the splits used in our experiments.
@@
We consider backbone instruction fine-tuned LLMs from different families of similar size. These are Llama-3.1-8B-Instruct (AI@Meta, 2024), Mistral-7B-Instruct-v0.3 Jiang et al. (2023), and Gemma2-9B-it Riviere et al. (2024). We also assess answer uncertainty quantification performance on QA models of the same family but with different sizes. To this end, we additionally evaluate on Gemma2-2B-it and Gemma2-27B-it. For all QA models, we use a simple prompt including the retrieved passages and the question in the context; the prompt is shown in Table 6 in the Appendix. The QA models’ response is the most likely answer generated with greedy sampling at temperature equal to 0. Following previous work on retrieval augmented QA, we use Contriever-MSMARCO Izacard et al. (2022) as our external retriever (Asai et al., 2024) and the target QA models are prompted with $|R|=5$ passages Yu et al. (2023); Asai et al. (2024); Xu et al. (2024). In Appendix B.2, we provide more details about inference and passage retrieval.
@@
A precise metric for measuring accuracy is key when evaluating the quality of uncertainty estimation. Token overlap metrics are imprecise and can over- or under-estimate accuracy (e.g., 5 will not match five). Thus, we rely on an LLM-based accuracy evaluator to create training data for the Passage Utility predictor (i.e., $A$ in Section 3.1) and to evaluate retrieval augmented QA performance. We use Qwen2-72B-Instruct (Yang et al., 2024) and the prompt proposed in Sun et al. (2024) to obtain accuracy judgments. More details about the LLM evaluator can be found in Appendix B.2.
@@
To assess the quality of answer uncertainty prediction, we follow Farquhar et al. (2024) and report the Area Under the Receiver Operator Curve (AUROC) on detecting incorrect answers (i.e., answers with high uncertainty). In Appendix C.3, we also report the area under the ‘rejection accuracy’ curve (AURAC) which captures the accuracy a model would have if it refused to answer questions with highest uncertainty. Rejection accuracy is essentially the model’s accuracy on the remaining questions. We report accuracy at different answer rejection thresholds, i.e., when models answer 80% and 90% of the least uncertain questions, as well as when always answering. We provide implementation details in Appendix B.2.
Training the Passage Utility Predictor
We train a Passage Utility predictor per QA model and QA task. For each task, we create set $D_{\mathcal{M}}=\{(x,p,\upsilon_{\mathcal{M}})\}$ to train and evaluate a Passage Utility predictor for QA model $\mathcal{M}$. We use the train (and development) questions available for the QA task and consider the top five retrieved passages for each question (i.e., $|R|=5$). Note that this is a hyper-parameter and other values would be also possible. Larger sizes of $|R|$ would yield more training data, since the Utility predictor takes individual passages (together with the question) as input, First, the target QA model $\mathcal{M}$ is prompted with passage $p\in R$ and question $x$ to generate answer $y$. Then, we annotate passages $p$ with a utility score computed with the accuracy evaluator $A$ and the entailment judge $E$ on the generated answer $y$ (Section 3.1). We use an ALBERT-xlarge Lan et al. (2020) model optimized on MNLI (Williams et al., 2018) and the VitaminC dataset (Schuster et al., 2021). We report further details about the Passage Utility predictor training in Appendix B.2.
@@
We compare against several strong uncertainty estimation methods (Fadeeva et al., 2023) which we briefly describe below and also report additional comparisons in Appendix C.3.
Information-Theoretic Measures We compare against uncertainty estimation methods that are based on the predictive probabilities of the target QA model. For a generated answer $y$ with probability $p(y|x,R;\mathcal{M})=\prod_{t=1}^{|y|}p(y_{t}|y_{1..t-1},x,R;\mathcal{M})$, the Perplexity (PPL) of model $\mathcal{M}$ is computed as:
$\text{PPL}(x,R,\mathcal{M})=$ (6)
$\exp{\{-\,\frac{1}{|y|}\,\sum^{|y|}_{t=1}\,\log\,p(y_{t}|y_{1..t-1},x,R;\mathcal{M})}\},$
Perplexity essentially calculates token-level entropy as it is based on the average negative log-likelihood of the generated tokens.
@@
$\text{RE}(x,R,\mathcal{M})=$ (7)
$-\frac{1}{N}\,\sum_{n=1}^{N}\text{log}\,\tilde{P}(y^{n}\,|\,x,R;\mathcal{M}),$
@@
$\text{SE}(x,R,\mathcal{M})=$ (8)
$-\sum^{M}_{m=1}\hat{P}_{m}(x,\mathcal{M})\,\text{log}\,\hat{P}_{m}(x,\mathcal{M}),$
where $\hat{P}_{m}(x,\mathcal{M})=\frac{\sum_{y\in C_{m}}\,\tilde{P}(y|x,R;\mathcal{M})}{\sum^{M}_{m=1}\,\sum_{y\in C_{m}}\,\tilde{P}(y|x,R;\mathcal{M})}$.
LLM-based Measures We compare with p(true) which uses the same LLM-based target QA model to assess whether the answers it produces are correct Kadavath et al. (2022). We follow the p(true) variant used in previous work (Kuhn et al., 2023). The QA model is prompted with the question and a set of candidate answers (consisting of the most likely answer and a sample of $N$ answers) and is instructed to respond whether the most likely answer is true or false (i.e., correct/incorrect). The score produced by this approach is the probability of model $\mathcal{M}$ generating the token True. This method needs several in-context examples to work well; following Kuhn et al. (2023), we use 20 in-context examples. Note that since our backbone LLMs are recent models with a large context (unlike Kuhn et al. 2023), all 20 examples are fed in the context making p(true) an expensive but very strong approach. In the context of retrieval augmented QA, we include in the prompt the set of retrieved passages for the question to evaluate. We illustrate the prompt used by p(true) in Appendix B.3.
@@ 5 Results and Analysis
Light-weight answer uncertainty prediction works across model families and question answering tasks.
Uncertainty estimation AUROC results for three QA models (Gemma2-9B, Llama-3.1-8B, and Mistral-7B-v0.3) are shown in Table 1 (results on the development set are included in Appendix C.3). In general, answer perplexity (PPL) performs rather poorly, especially for Gemma2-9B and Mistral-7B-v0.3. Regular Entropy improves upon PPL by ignoring surface form choices and focusing on meaning, Semantic Entropy further improves AUROC scores. p(true) performs well at detecting answer uncertainty matching or surpassing Semantic Entropy. It is important to note that this method relies on sampled answers and a long prompt with 20 in-context examples. Our Passage Utility approach performs on par or outperforms all other methods with a single inference step on each input passage.
Passage Utility performs particularly well on challenging question answering tasks represented by datasets like WebQ, SQuAD, and RefuNQ. In these cases, our light-weight uncertainty estimation model works better than p(true) which requires the same QA model (i.e., the same backbone LLM) to judge the correctness of its own generated answers. We speculate that for questions with high uncertainty, i.e., where the model does not have the knowledge to answer, it confidently generates a response and also fails at assessing it (e.g., questions about non-existing concepts in RefuNQ). We attribute the Passage Utility’s success to the fact that it has been specifically trained to detect situations where the target QA model is prone to answer incorrectly (i.e., when provided with retrieved passages of lower relevance).
NQ TQA WebQ SQuAD PopQA RefuNQ
Gemma2-9B
PPL 0.64 0.68 0.52 0.53 0.59 0.51
p(true) 0.73 0.75 0.67 0.63 0.81 0.62
Regular Entropy 0.66 0.69 0.54 0.56 0.61 0.51
Semantic Entropy 0.70 0.73 0.57 0.64 0.73 0.59
Passage Utility 0.72 0.82 0.70 0.79 0.84 0.81
Llama-3.1-8B
PPL 0.75 0.80 0.68 0.74 0.83 0.60
p(true) 0.79 0.88 0.72 0.77 0.85 0.67
Regular Entropy 0.76 0.81 0.69 0.78 0.83 0.65
Semantic Entropy 0.71 0.78 0.69 0.78 0.79 0.58
Passage Utility 0.78 0.78 0.76 0.82 0.86 0.82
Mistral-7B-v0.3
PPL 0.63 0.71 0.57 0.65 0.64 0.62
p(true) 0.73 0.82 0.68 0.74 0.75 0.68
Regular Entropy 0.64 0.75 0.62 0.65 0.66 0.60
Semantic Entropy 0.66 0.77 0.66 0.74 0.74 0.61
Passage Utility 0.74 0.82 0.70 0.81 0.86 0.80
Table 1: AUROC values for QA models Gemma2-9B, Llama-3.1-8B, and Mistral-7B-v0.3 on Natural Questions (NQ), TriviaQA (TQA), WebQuestions (WebQ), SQuAD, PopQA, and RefuNQ test sets. Best values are highlighted in bold.
0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.8Gemma2-9B696969696767676766666666727272726969696966666666707070706868686866666666727272726969696966666666747474747070707066666666AccLMPPLRegular EntropyUtility Rankerp(true)Semantic Entropy
0.50.50.50.50.60.60.60.60.70.70.70.7Llama-3.1-8B707070706666666662626262707070706666666662626262717171716767676762626262707070706666666662626262717171716767676762626262AccLM
80% 90% 100% 0.50.50.50.50.60.60.60.60.70.70.70.7Mistral-7B-v0.3666666666565656563636363727272726868686863636363686868686666666663636363707070706767676763636363727272726868686863636363AccLM
Figure 2: Average QA model accuracy (AccLM) across test sets: models refuse to answer according to different uncertainty estimation methods. 80/90%: the model answers 80/90% of the questions with low uncertainty; 100%: the model answers all questions.
Answering selectively based on Passage Utility improves QA accuracy.
Answer uncertainty can be used to decide whether to answer or refrain from doing so. Figure 2 shows QA model accuracy across datasets (average AccLM) at different thresholds of answer rejection. We report accuracy when models choose to answer only 80% and 90% of the cases where they are most certain, and when they always answer. Across different LLM backbones, Passage Utility performs on par with or better than more expensive uncertainty estimation approaches.
@@
Passage Utility performance remains consistent across model sizes.
@@
Figure 3 (left), shows average AUROC scores on answer uncertainty estimation for Passage Utility and comparison approaches with different Gemma2 sizes: 2B, 9B, and 27B. Our Passage Utility approach performs best across model sizes. All approaches obtain better AUROC scores for the smaller Gemma2 model (2B) which makes most errors; we observe a noticeable decrease in performance for most information-theoretic models when using the biggest Gemma model (27B). We attribute this to the fact that the 27B model more confidently makes less errors. p(true) on the other hand benefits from the largest model’s context understanding and memorised knowledge.
Figure 3 (right) shows average accuracy when the target QA models choose to answer 80% of the cases they are most confident about. For comparison, we also show QA accuracy when always answering, i.e., black bold dots. When looking at selective performance according to Passage Utility, the small Gemma2-2B model surpasses the bigger Gemma2-27B one when always answering (0.68 vs 0.65), whereas the biggest Gemma2-27B model improves by +9 points (0.74 vs 0.65).
@@
2B9B27B0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.8Avg AUROCPPLRegular EntropyPassage Utilityp(true)Semantic Entropy
2B9B27B0.60.60.60.60.650.650.650.650.70.70.70.70.750.750.750.75Avg AccLM on 80%
Figure 3: (left) Average AUROC for answer uncertainty estimation methods with varying Gemma2 sizes: 2B, 9B, and 27B. (right) Average AccLM on the 80% most confident answers; black dots indicate average AccLM when always answering. Results computed on test sets.
Passage Utility scores provide good passage re-ranking.
We hypothesize that by re-ordering the set of retrieved passages according to their Passage Utility score and filtering the top-$k$ ones, it will be possible to improve the accuracy and efficiency of the target QA model Salemi and Zamani (2024); Liu et al. (2024b). To test this, we measure QA accuracy for Gemma2-9B on sets of input passages varying in order and size. Specifically, we measure performance with a set of $|R|=10$ passages in their original ranking provided by an external retriever and the re-ranking imposed by the Passage Utility score. We then compute accuracy for the top-$k$ passages with $k$ in the range of {10, 5, 3, 1}. Figure 4 shows average accuracy (AccLM) values across five datasets (NQ, TQA, WebQ, SQuAD, and PopQA) on different input sets created based on the original retriever (gray) and the Passage Utility scores (red). At $k=10$ both rankings lead to the same average QA accuracy (67%). However, when reducing the size of the context to the top $k=5,3,1$ passages re-ranked by the Passage Utility score, the QA model achieves higher accuracy, indicating that Passage Utility indeed captures which passages are useful for the target QA model.
105310.550.550.550.550.60.60.60.60.650.650.650.65Avg AccLMRetriever ScorePassage Utility
@@
Figure 4: Average retrieval augmented QA accuracy (AccLM) for Gemma2-9B across five QA test sets (NQ, TQA, WebQ, SQuAD, and PopQA). Points on the x-axis correspond to different context sizes $|R|$, when taking the top-$k$ passages from different rankings produced by the retriever and Passage Utility estimator; $k$ varies over {10, 5, 3, 1}.
Training with pairwise judgements on Passage Utility helps improve predictions.
Table 2 shows AUROC results on answer uncertainty prediction with Passage Utility estimators trained with different variants of the objective in Equation (4). The first row shows the full objective (see training details in Section B.2), the second row shows a variant where the ranking objective uses only entailment annotations ($e$), and in the third row the objective is solely based on accuracy prediction ($\mathcal{L}_{BCE}$). As can be seen, there is a drop in performance when the pairwise ranking loss is not used (i.e., last line of Table 2); this component of the objective provides a smoother signal on passage utility which is empirically beneficial. However, when the pairwise ranking loss is only based on entailment (an external critic), performance drops by several points which highlights the importance of training with utility judgements provided by the target QA model.
G9B L8B M7B
$\mathcal{L}_{rank},(e+a)/2+\lambda\,\mathcal{L}_{BCE}$ 0.79 0.81 0.79
$\mathcal{L}_{rank},(e)+\lambda\,\mathcal{L}_{BCE}$ 0.71 0.73 0.71
$\mathcal{L}_{BCE}$ 0.77 0.78 0.80
Table 2: Answer uncertainty estimation with Passage Utility predicted by models trained with different variants of the training objective in Equation 4. We report AUROC for Gemma2-9B (G9B), Llama3.1-8B (L8B), and Mistral-7B-v0.3 (M7B) averaged over development sets.

### v2

@@ 3.2 Answer Uncertainty Estimation
For our QA task, we want to define an estimator $\{x,R\}\mapsto\text{{u}}_{\mathcal{M}}(\{x,R\})$ which quantifies the uncertainty of model $\mathcal{M}$ when generating answer $y$ for question $x$ based on a prompt with passages $R$. We propose estimating $\text{{u}}_{\mathcal{M}}$ directly from the utility scores of individual passages in $R$. The key intuition is that the higher the utility of one (or more) passages, the less uncertain $\mathcal{M}$ will be when generating answer $y$. Conversely, if all passage utilities in $R$ are low, it is more likely that $\mathcal{M}$ will be uncertain about the answer. Specifically, we estimate answer uncertainty $\textbf{u}_{\mathcal{M}}$ by taking the maximum utility score among the passages in $R$:
@@
Our approach to aggregating passage utilities is intuitive and simple. However, QA models might be sensitive to factors relating to how they are prompted, such as the number or order of passages (Liu et al., 2024b; Xie et al., 2024). In Appendix C, we examine such confounds more closely, comparing question answering accuracy when models are prompted with individual passages in $R$ (our aggregation approach) versus the entire set $R$. The study shows that there is little disagreement between the two methods and that it is possible to approximate answer uncertainty when prompting with $|R|$ passages while avoiding the combinatorial complexity of estimating uncertainty over all possible combinations of input passages.
@@ 4 Experimental Setup
We evaluate our approach to predicting answer uncertainty on short-form question answering tasks. Specifically, on the following six datasets: Natural Questions (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), WebQuestions (Berant et al., 2013), SQuAD (Rajpurkar et al., 2016), and PopQA (Mallen et al., 2023). We also evaluate on RefuNQ (Liu et al., 2024a), a dataset with unanswerable questions about non-existing entities. In Appendix E.1, we describe each dataset, provide example questions, and make available details about the splits used in our experiments which follow Lee et al. (2019).
@@
We consider backbone instruction fine-tuned LLMs from different families of similar size. These are Llama-3.1-8B-Instruct (AI@Meta, 2024), Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), and Gemma2-9B-it (Riviere et al., 2024). We also assess answer uncertainty quantification performance on QA models of the same family but with different sizes. To this end, we additionally evaluate on Gemma2-27B-it. For all QA models, we use a simple prompt including the retrieved passages and the question in the context; the prompt is shown in Table 9 in the Appendix. The QA models’ answer is the most likely answer generated with greedy sampling at temperature equal to 0. Following previous work on retrieval augmented QA, we use Contriever-MSMARCO (Izacard et al., 2022) as our external retriever (Asai et al., 2024) and the target QA models are prompted with $|R|=5$ passages (Yu et al., 2023; Asai et al., 2024; Xu et al., 2024). In Appendix E.2, we provide more details about inference and passage retrieval.
@@
A precise metric for measuring accuracy is key when evaluating the quality of uncertainty estimation. Token overlap metrics are imprecise and can over- or under-estimate accuracy (e.g., 5 will not match five). Thus, we rely on a LLM-based accuracy evaluator to create training data for the Passage Utility predictor (i.e., $A$ in Section 3.1) and to evaluate retrieval augmented QA performance. We use Qwen2-72B-Instruct (Yang et al., 2024) and the prompt proposed by Sun et al. (2024) to obtain accuracy judgments. Details about the LLM evaluator can be found in Appendix E.2.
@@
To assess the quality of answer uncertainty prediction, we follow Farquhar et al. (2024) and report the Area Under the Receiver Operator Curve (AUROC) on detecting incorrect answers (i.e., answers with high uncertainty). In Appendix F.2, we also report the area under the ‘rejection accuracy’ curve (AURAC) which captures the accuracy a model would have if it refused to answer questions with highest uncertainty. Rejection accuracy is essentially the model’s accuracy on the remaining questions. In the main results section, we focus on selective answering performance when models answer 80% of the least uncertain questions versus when always answering. We provide implementation details in Appendix E.2.
Training of Passage Utility Predictor
We train a Passage Utility predictor per QA model and QA task. For each task, we create set $D_{\mathcal{M}}=\{(x,p,\upsilon_{\mathcal{M}})\}$ to train and evaluate a Passage Utility predictor for QA model $\mathcal{M}$. We use the training (and development) questions available for each QA task, considering the top five retrieved passages for each question (i.e., $|R|=5$). Note that $|R|$ a hyper-parameter and other values would be also possible. Larger sizes of $|R|$ would yield more training data, since the Utility predictor takes individual passages (together with the question) as input, The target QA model $\mathcal{M}$ is first prompted with passage $p\in R$ and question $x$ to generate answer $y$. Then, we annotate passages $p$ with a utility score computed with the accuracy evaluator $A$ and entailment judge $E$ on generated answer $y$ (Section 3.1). We use an ALBERT-xlarge model (Lan et al., 2020) optimized on MNLI (Williams et al., 2018) and the VitaminC dataset (Schuster et al., 2021). We report further details about the Passage Utility predictor training in Appendix E.2.
@@
There exist several uncertainty estimation methods which we group in two categories based on whether they require access to logits or simply model outputs (see Fadeeva et al. 2023 for additional methods). We choose the highest scoring ones to compare with here and include additional results in Appendix F.2 for completeness.
Information-Theoretic Measures. We compare against uncertainty estimation methods that are based on the predictive probabilities of the target QA model. Let $y$ denote an answer generated with probability $p(y|x,R;\mathcal{M})$ which is computed as:
$p(y|x,R;\mathcal{M})=\prod_{t=1}^{|y|}p(y_{t}|y_{1..t-1},x,R;\mathcal{M})$ (7)
The Perplexity (PPL) of model $\mathcal{M}$ boils down to calculating token-level entropy as it is based on the average negative log-likelihood of the generated tokens:
$\text{PPL}(x,R,\mathcal{M})=\exp{\{-\,\frac{1}{|y|}\,\sum^{|y|}_{t=1}\,\log\,p(y_{t}|y_{1..t-1},x,R;\mathcal{M})}\},$ (8)
@@
$\text{RE}(x,R,\mathcal{M})=-\frac{1}{N}\,\sum_{n=1}^{N}\text{log}\,\tilde{P}(y^{n}\,|\,x,R;\mathcal{M}),$ (9)
@@
$\text{SE}(x,R,\mathcal{M})=-\sum^{M}_{m=1}\hat{P}_{m}(x,\mathcal{M})\,\text{log}\,\hat{P}_{m}(x,\mathcal{M}),$ (10)
where $\hat{P}_{m}(x,\mathcal{M})$ is estimated as follows:
$\hat{P}_{m}(x,\mathcal{M})=\frac{\sum_{y\in C_{m}}\,\tilde{P}(y|x,R;\mathcal{M})}{\sum^{M}_{m=1}\,\sum_{y\in C_{m}}\,\tilde{P}(y|x,R;\mathcal{M})}$ (11)
LLM-based Measures. We compare with p(true) which uses the same LLM-based target QA model to assess whether the answers it produces are correct (Kadavath et al., 2022). We follow the p(true) variant used in previous work (Kuhn et al., 2023). The QA model is prompted with the question and a set of candidate answers (consisting of the most likely answer and a sample of $N$ answers) and is instructed to respond whether the most likely answer is true or false (i.e., correct/incorrect). The score produced by this approach is the probability of model $\mathcal{M}$ generating the token True. This method needs several in-context examples to work well; following Kuhn et al. (2023), we use 20 in-context examples. Note that since our backbone LLMs are recent models with a large context (unlike Kuhn et al. 2023), all 20 examples are fed in the context making p(true) an expensive but very strong approach. In the context of retrieval augmented QA, we modify p(true) to include in the prompt the set of retrieved passages for the question of interest. We provide the prompt used by p(true) in Appendix E.3. Note that p(true) can be considered as a specialized powerful entailment model and thus we do not include entailment based methods relying on off-the-shelf NLI models which have been shown to perform poorly (Yoran et al., 2024).
@@ 5 Results and Analysis
Passage Utility is effective across model families, sizes, and QA tasks.
Table 1 summarizes our uncertainty estimation results (test set) with four QA models (Gemma2-9B/27B, Llama-3.1-8B, and Mistral-7B-v0.3) across six QA datasets (results on the development set are included in Appendix F.2). We boldface the highest AUROC value for each QA model and dataset pair and mark with  the next best value that is significantly different from it at $p<0.05$. We use the paired De Long test (DeLong et al., 1988) to compute whether two AUROC values are significantly different. [^2] We use the library in https://github.com/Brritany/MLstatkit to compute significance scores.
In general, answer perplexity (PPL) performs rather poorly, especially with Gemma2-9B/27B. Perplexity is likely to underperform with less calibrated models, such as those which have undergone instruction tuning (Tian et al., 2023). Regular Entropy shows little improvement upon PPL but by ignoring surface form choices and focusing on meaning, Semantic Entropy improves AUROC scores. p(true) performs well at detecting answer uncertainty matching or surpassing Semantic Entropy. Overall, we observe that the gap among these methods’ performance is lower than in the context of closed-book QA studied in previous work (Farquhar et al., 2024). We hypothesize that, on one hand, our recent QA models admitting more in-context examples benefiting p(true) and, on the other hand, that retrieved passages in the prompt make QA models’ outputs less varied. Our Passage Utility approach performs on par or outperforms all other methods with a single small-model inference step on each input passage.
Passage Utility performs particularly well on challenging question answering tasks represented by datasets like PopQA and RefuNQ. In these cases, our light-weight uncertainty estimation model works better than p(true) which requires the same QA model (i.e., the same backbone LLM) to judge the correctness of its own generated answers. We speculate that for questions with high uncertainty, i.e., where the model does not have the knowledge to answer (e.g., questions about non-existing concepts in RefuNQ), it confidently generates a response and also fails at assessing it. We attribute the Passage Utility’s success to the fact that it has been specifically trained to detect situations where the target QA model is prone to answer incorrectly (i.e., when provided with retrieved passages of lower relevance). The six QA tasks pose different retrieval challenges. On TQA, retrieval results are often of good quality: for 73% of the questions, the top-5 retrieved passages contain the gold answer string. In contrast, on PopQA the percentage reduces to 63% and on RefuNQ the quality of retrieval is deliberately low (as it consists of unanswerable questions). Across models (Table 1), our approach is comparable to p(true) when retrieved passages contain the answer and excels in cases of low quality retrieval.
Gemma2-9B Gemma2-27B
NQ TQA WebQ SQuAD PopQA RefuNQ AVG NQ TQA WebQ SQuAD PopQA RefuNQ AVG
PPL 0.64 0.68 0.52 0.53 0.59 0.51 0.58 0.64 0.50 0.53 0.59 0.58 0.51 0.56
p(true) 0.73   0.75 0.67 0.63   0.81 0.62 0.70 0.77 0.83 0.67   0.68   0.78   0.60 0.72
Regular Entropy 0.66 0.69 0.54 0.56 0.61 0.51 0.60 0.67 0.54 0.55 0.59 0.62 0.51 0.58
Semantic Entropy 0.70 0.73   0.57   0.64 0.73   0.59 0.66 0.69   0.62   0.59 0.63 0.66 0.58 0.63
Utility Ranker 0.76 0.85 0.69 0.78 0.86 0.79 0.79   0.73 0.82 0.69 0.80 0.85 0.78 0.78
Llama-3.1-8B Mistral-7B-v0.3
NQ TQA WebQ SQuAD PopQA RefuNQ AVG NQ TQA WebQ SQuAD PopQA RefuNQ AVG
PPL 0.75 0.80 0.68 0.74 0.83 0.60 0.73 0.63 0.71 0.57 0.65 0.64 0.62 0.64
p(true) 0.79 0.88 0.74 0.77 0.85   0.67 0.78 0.73 0.82 0.68   0.74    0.75   0.68 0.73
Regular Entropy   0.76 0.81   0.71 0.78   0.83 0.65 0.76 0.64 0.75   0.62 0.65 0.66 0.60 0.65
Semantic Entropy 0.72   0.82 0.66   0.78 0.81 0.59 0.73   0.66   0.78 0.66 0.73 0.74 0.61 0.70
Utility Ranker 0.77 0.82 0.72 0.83 0.87 0.81 0.80 0.74 0.83 0.68 0.82 0.85 0.80 0.79
Table 1: AUROC values for QA models based on Gemma2-9B/27B, Llama-3.1-8B, and Mistral-7B-v0.3 on Natural Questions (NQ), TriviaQA (TQA), WebQuestions (WebQ), SQuAD, PopQA, and RefuNQ test sets. The best values (per model and dataset) are highlighted in bold; we also mark with next best values which are significantly different using the paired De Long test ($p<0.05$). For example, on TQA with Gemma2-9B, p(true), the second best performing is significantly different from the Utility Ranker which performs best and by extension models with lesser values than p(true) are also significantly different.
@@
Passage Utility also performs well with different QA model sizes (within the same family), i.e., Gemma2-9B and 27B (Table 1). We observe a noticeable decrease in performance for most information-theoretic models when using the biggest Gemma model (27B). We attribute this to the fact that the 27B model more confidently makes less errors and its calibration may be affected more by the fine-tuning step (Tian et al., 2023). p(true), on the other hand, benefits from the largest model’s context understanding and memorized knowledge.
It is important to note that our approach is lite-weight at inference time. In Figure 2 we report average AUROC per method with respect to the number of input tokens at inference time and the number of parameters involved in uncertainty estimation. We report scores for Gemma2-9B and its bigger version Gemma2-27B. As can be seen, our approach which is based on a BERT encoder (Devlin et al., 2019) with 110M parameters and a small number of input tokens achieves on aggregate better performance than more expensive approaches. For the 27B QA model, p(true) edges closer to passage utility, however, at the expense of thousands of input tokens and $\sim$26 billions more parameters. This indicates that p(true) will be less efficient in QA settings where latency and cost are critical. In Appendix B we provide a cost analysis in terms of model calls.
7009k13k0.50.50.50.50.60.60.60.60.70.70.70.70.80.80.80.8PPL (27B)p(true) (27B)SE (27B)PU (27B)PPL (9B)p(true) (9B)SE (9B)PU (9B)Number of Inference TokensAverage AUROCGemma2-9BGemma2-27B
Figure 2: Average AUROC across our six development sets (y-axis) with respect to number of input tokens at inference time (x-axis) and number of parameters (size of the circles). We present results for perplexity (PPL), Semantic Entropy (SE), p(true), and Passage Utility (PU). We exclude Entropy (which is close to SE) for readability. We compare the smaller Gemma2-9B and its bigger version 27B. Thinner circles positioned in the left corner are better.
@@
G9BG27BL8BM7B0.60.60.60.60.650.650.650.650.70.70.70.70.750.750.750.75Model variantsAverage QA AccuracyPPLRegular EntropyPassage Utilityp(true)Semantic Entropy
Figure 3: Average accuracy (across our six test sets) with Gemma2 sizes 9B and 27B (G9B, G27B), Llama-3.1-8B (L8B), and Mistral-7B-v0.3 (M7). Black dots: QA models always answer; colour dots: QA models answer 80% of the cases they are most confident about.
Passage Utility leads to selective answering.
Model uncertainty can be used to decide whether to provide an answer to a question or not. Figure 3 shows average accuracy when the target QA models choose to answer 80% of the cases they are most confident about. For comparison, we also show QA accuracy when always answering, i.e., black bold dots. All uncertainty quantification approaches improve model accuracy. Across different LLM QA backbones, Passage Utility performs on par with or better than more expensive uncertainty estimation approaches such as p(true). For instance, when looking at selective performance according to Passage Utility, the biggest Gemma2-27B model improves by +9 points (0.74 vs 0.65). In Appendix F.2 we report the full set of selective accuracies at different thresholds.
Passage Utility shows potential for retrieval reranking.
To further assess the quality of the Passage Utility scores and to highlight their potential for retrieval reranking we carry out the following ablation study. Since most previous work (Asai et al., 2024; Xie et al., 2024; Yoran et al., 2024) on retrieval augmented QA prompts the QA model with the top-$5$ (or less) input passages, we hypothesize that our passage utility score could be an effective reranking method after retrieval (Nogueira et al., 2019; Ma et al., 2024; Yao et al., 2024). We test this hypothesis by computing the accuracy of the Gemma2-9B QA model when prompted with the top-$k$ passages (with $k$ in the range of {5, 3, 1}) out of a sample of $|R|=10$ passages provided by an external retriever.
In Figure 4, we compare performance under two passage-ranking strategies: one based on relevance scores from an external retriever (gray), and the other based on the QA model’s self-assessed utility of individual passages, following Yao et al. (2024). We report two self-assessment variants, one using the predictions of the Passage Utility model (red) and another one based on the perplexity of the QA model when answering with individual retrieved passages (blue). Passages that yield answers with lower perplexity should be ranked first. Figure 4 shows average accuracy values across five development sets (NQ, TQA, WebQ, SQuAD, and PopQA), at different cutoff values ($k$). As can be seen, the QA model achieves higher accuracy when passages are ranked according to their utility. This finding suggests that Passage Utility scores indeed reflect which passages are useful for the target QA model.
Passage Utility generalizes to complex reasoning questions.
We assess the generalization of our approach on multi-hop questions, i.e., questions that require reasoning over pieces of information distributed across more than one passage (Yang et al., 2018; Pal et al., 2022). We evaluate on the widely used HotPotQA dataset (Yang et al., 2018), we use the splits as provided by Trivedi et al. (2023), and the same retrieval settings as defined above (Section 4). Table 2 shows AUROC values for all uncertainty estimation methods.
Across models, the performance of our approach is on par with p(true) and sequence entropy methods, and better than perplexity. Manual inspection of 100 examples from the development set with Gemma2-9B as the target QA model, reveals two major trends. Firstly, QA models can perform the multi-hop task by relying on memorized knowledge and the provided passages. In 48 cases, the QA model can correctly answer with several of the passages in $R$ when prompted individually; the model also correctly answers when prompted with $R$. Secondly, the retrieved passages tend to be of low quality. For 29 examples, the model cannot answer neither when prompted with individual passages nor with the entire set $R$. Our approach works well in these two cases with few false positives (negatives). Beyond the assessment of answer uncertainty estimation, this evaluation shows that the idealized multi-hop behavior studied with synthetic passages (Yang et al., 2018) occurs to a lesser extent in current LLMs in the context of open book QA where it is more challenging to retrieve relevant passages in the first place. Note that multi-hop QA often requires more complex QA pipelines such as interactive retrieval (Trivedi et al., 2023).
5310.50.50.50.50.60.60.60.60.70.70.70.7Number of Retrieved PassagesAverage AccuracyRAGRAG+PURAG+PPL
@@
Figure 4: Average RAG accuracy with Gemma2-9B across the five QA development sets. Points on the x-axis correspond to different context sizes, when taking the top-$k$, passages according to query relevance (gray) and self-assessment via perplexity (PPL; blue), and Passage Utility (PU; red).
G9B G27B L8B M7B
PPL 0.61 0.59 0.73 0.70
p(true)   0.75 0.78 0.81 0.78
Regular Entropy 0.64 0.60   0.77 0.71
Semantic Entropy 0.70   0.65 0.74 0.73
Utility Ranker 0.78 0.77   0.77   0.74
Table 2: AUROC values for QA models based on Gemma2-9B/27B, Llama-3.1-8B, and Mistral-7B-v0.3 on HotPotQA test set. Best values per model are highlighted in bold; we also mark with next best values which are significantly different using the paired De Long test ($p<0.05$).